# State-Space Modeling in Long Sequence Processing: A Survey on Recurrence in the Transformer Era

### v1

@@ 7.3 Natural Language Benchmarks
Language Modeling.   One of the peculiar characteristic behind the success of LLMs is their ability to model the generative likelihood of word sequences, i.e., they are able to predict the probabilities of future tokens. In other words, they can build powerful Language Models for a target language, being able to capture the syntax, grammar, and semantic properties. There exist several Language Modeling dataset, and many of the works described in this survey are evaluated in Language Modeling tasks. PennTreebank (PTB) [181], WikiText-103 [192] and The Pile [72] are commonly used for evaluating model performances terms of perplexity. PTB is a language model corpus containing 1 million words. Generally, models are trained to predict the next word (a variant denoted with PTB-w, evaluated with the perplexity metric) or character (referred to as PTB-c, evaluated with bits-per-character metric). The corpus contains $42,068$ sentences ($971,657$ words, average word-length of about $4.727$ characters) of varying length (the range is from $3$ to $84$ words). The vocabulary for the character level models includes $49$ unique symbols (including one for spaces). The standard train/valid/test split consists of $929$k training words, $73$k validation words, and $82$k test words, with lower-case words, numbers replaced with token $N$, newlines replaced with $<$eos$>$, all other punctuation removed. The vocabulary is composed by the most frequent $10$k words, with the rest of the tokens replaced by an $<$unk$>$ token [196]. In literature, various segmentations of the corpus into training sequences of different lengths are being used ($L=50$ for PTB-w in [157, 250], $L\in(70,300)$ for PTB-w and $L=150$ for PTB-c in [130]). This task is considered in [157, 130, 51, 25, 140, 131, 250, 209]. WikiText-103 [192] is a long-range world-level language modeling benchmark. It contains $28$k articles from English Wikipedia, with an average length of $3.6$k tokens per article. In total, it contains $103$M words. The preprocessing is more realistic with respect to PTB (e.g., vocabulary is not limited to $10$k words, only words with a count below $3$ are discarded). It is used in [89, 88, 171, 159, 227, 224, 118, 259, 230, 111, 51, 30, 158, 133, 180, 233, 304, 135, 235, 191]. The Pile dataset [72] is an $825$ GiB English text corpus designed for training large-scale language models. It is is composed of $22$ diverse and high-quality datasets, including established Natural Language Processing datasets (PG-19, Books3, English Wikipedia, etc.) and $14$ new high-quality curated language modeling datasets derived from heterogeneous sources (PubMed, ArXiv, GitHub, Stack Exchange, etc.–see Table 1 in [72] for a full description)). Besides its impact in training large language models, The Pile can serve as a benchmark for testing cross-domain knowledge and generalization ability of language models. It is exploited in [227, 85, 277, 29, 235, 222]. Among the datasets contained in The Pile, the Project Gutenberg (PG-19) [238] consists of $28,602$ full-length books from the Project Gutenberg, published prior to $1919$, hence representing distinct writing styles with respect to other datasets based on modern text. It contains $6,966,499$ English language words. When tokenized, each PG-19 book has between $50$k-$100$k tokens. This datasets has been employed to test long range modeling capabilities of the approaches proposed in [225, 114, 304]. Datasets from Arxiv and GitHub were already been introduced in [305], prior to The Pile. The ArXiv dataset is a corpus of sources of technical papers from the “Mathematics” section, downloaded via the ArXiv Bulk Data Access. The GitHub dataset consists of source code extracted from diverse GitHub repositories featuring open-source licenses. In this dataset, all files within each GitHub repository are combined to form one long document. The evaluation metric commonly used with these datasets is bits-per-token (i.e., $\text{log}_{2}(\text{perplexity})$, the lower the better). Used by [114, 225, 158]. Enwik8 [176] is a character-level dataset consisting of $100$M unprocessed bytes of the XML text dump of the English Wikipedia. Its vocabulary contains $204$ characters, and it is exploited in [313, 111, 51, 30, 153].
@@
Commonsense Reasoning.   A common experimental setting for evaluating LLMs consists in pre-training the model on The Pile and then testing the model capabilities on downstream zero-shot evaluation tasks. In particular, recent works leverage the Language Model Evaluation Harness (LM-Eval-Harness) [^24] https://github.com/EleutherAI/lm-evaluation-harness from EleutherAI [73], which mainly evaluates on tasks/datasets that measure common sense reasoning—in which the model should use “common sense” or world knowledge to perform inference. It is composed by several tasks: LAMBADA, HellaSwag, PIQA, ARC, WinoGrande. LAMBADA [214] tests the modeling capacity on long-range contextual reasoning and language comprehension abilities. The model is given a certain context sentence, and it is asked to predict the last word of a target sentence based on the context. Accuracy and perplexity of the predicted last words are measured to assess the performance on language modeling tasks. The scaling law property [132, 106] is often measured, meaning that scaling language models would improve the accuracy and reduce the perplexity. LAMBADA contains novels from the Book Corpus [317], and comprises $10,022$ passages, divided into $4,869$ development (that can be used to finetune models) and $5,153$ test passages (extracted from $1,331$ and $1,332$ disjoint novels, respectively). The average passage consists of $4.6$ sentences in the context plus $1$ target sentence, for a total length of $\approx 75$ tokens. The dataset provides also training data, which include the full text of $2,662$ novels (disjoint from those in dev/test sets), comprising $203$ million words. HellaSwag [312] [^25] Short for Harder Endings, Longer contexts, and Lowshot Activities for Situations With Adversarial Generations. expands the SWAG dataset [311], which introduced the task of commonsense natural language inference: the model receives context from a video caption, along with four potential outcomes for the subsequent events. Among the provided choices, only one is correct, representing the actual succeeding caption in the video. Negatives are created using Adversarial Filtering, a method where a set of discriminators is employed to choose a tricky collection of incorrectly generated answers. According to the HellaSwag paper, models like BERT struggle with robust commonsense reasoning and tend to learn specific biases from the dataset. If there’s a slight shift in language distribution, their performance drops significantly, even if the domain remains the same. HellaSwag introduces $70,000$ problems that are easy for humans ($95.6$% accuracy) but challenging for machines ($50$%). These problems consist of video captions from the ActivityNet Captions dataset and context and follow-up paragraphs from WikiHow. To assess how a model can adapt to novel scenarios, category labels sourced from WikiHow and ActivityNet are sampled to build ’zero-shot’ evaluation sets. For each set, being it used for validation or testing, two subsets are generated. The first one includes $5,000$ ’in-domain’ examples from categories encountered during training. The second subset comprises $5,000$ ’zero-shot’ examples randomly selected from categories held out during training. In total, there are $70,000$ examples in the dataset. PIQA (Physical Interaction: Question Answering) [22] benchmarks are about physical commonsense understanding. Given a “physical” goal (“to separate egg whites using a water bottle…”) expressed in natural language and two possible solutions (($i$) “Squeeze the water bottle and press it against the yolk. Release, which creates suction and lifts the yolk;” ($ii$) “Place the water bottle and press it against the yolk. Keep pushing, which creates suction and lifts the yolk.”), a model must choose the most appropriate solution. The dataset assesses the capacity of natural language understanding models to connect text with a robust intuitive-physics model of the world. Humans effortlessly choose answer ($i$) because separating the egg involves pulling out the yolk, a task that machines find easily misleading. The dataset consists of more than $16,000$ training QA pairs with an additional $\approx 2$k and $\approx 3$k QA held out for development and testing, respectively. Goal sentences have an average length of $7.8$ words, while correct/incorrect solutions have an average length of $21.3$ words. This results in over $3.7$ million lexical tokens within the training data. The AI2 Reasoning Challenge (ARC) [46] consists of a collection of $7787$ natural grade-school science questions ( typically $4$-way multiple choice), authored for human standardized tests. It is partitioned into two sets: Challenge Set (ARC-c), which contains $2590$ questions which are incorrectly answered by both a retrieval-based algorithm and a word co-occurrence algorithm. An Easy Set (ARC-easy) contrains the remainder of the questions. WinoGrande [252] is a large-scale dataset of $44$k pronoun resolution problems. The task is to identify to which of two subject a pronoun [^26] In practice it is formatted as a fill-in-the-blank problem, where the blank corresponds to the mention of one of the two names in the context. is referring to (e.g., “Robert woke up at 9:00am while Samuel woke up at 6:00am, so _ had less time to get ready for school.”; Options: Robert / Samuel). Generally, there is a twin sentence with nearly identical structure but where the answer is the opposite. It has been designed to be unsolvable for statistical models that rely on basic word associations. Going back to the main container of the just described datasets, the LM-Eval-Harness, we mentioned that it has been used in several recent works [85, 57, 309, 222, 277, 231]. Apart from LM-Eval-Harness, other comprehensive benchmarks for commons sense reasoning have been tested. Measuring Massive Multitask Language Understanding (MMLU) [102] is a versatile benchmark for large-scale evaluation of multi-task knowledge understanding. It consists of $57$ tasks including elementary mathematics, US history, computer science, law, etc. It collects $15,908$ questions in total, split into a few-shot development set, a validation set, and a test set. As shown in existing work, LLMs mostly outperform small models by a substantial margin on this benchmark, also showing scaling laws in model size. We point the reader attention towards a recent survey on LLMs [316] for further details on other commonly used datasets, such as CMMLU [155], C-Eval [112] BoolQ [45], OpenBookQA [195] and others.
@@
Machine Translation.   Another sequence-oriented benchmark in the context of language generation is conditional text generation [156], which focuses on generating texts satisfying specific task demands based on the given conditions. A common task is machine translation [12], whose goal is the translation of text or speech from a source language to another. Among the most common benchmarks, the WMT (several editions—2014 [24] up to 2022 [146]) contains translation tasks from/to English to/from several languages, such as Czech, French, German, Hindi, and Russian. A common testbed is to use the WMT’14 English-to-German, which contains 4.6M sentence pairs, and the WMT’17 the Chinese-to-English split which is composed by $20.6$M sentence pairs. The case-sensitive NIST BLEU score [215] is commonly adopted as the evaluation metric. It has been tackled in [171, 224, 259, 97, 35, 133, 37]. A similar benchmark is the IWSLT German–English spoken-domain translation [33]. It consists of $209,772$ sentence pairs from transcribed TED and TEDx presentations, with a mean sentence length of $103$ characters for German and $93$ for English. It has been considered in [224, 25].
@@
Other Language-oriented Benchmarks.   Apart from the generation of high-quality natural language text, several models show strong abilities to generate regular and formal languages, such as computer programs (i.e., code synthesis) [111, 316]. Another popular direction for evaluation is the one is sentiment classification. The principal testbed is the IMDb movie review dataset [172], a collection of $50$k movie reviews, equally balanced into positive and negative reviews and processed into equal-size train and test sets. Sequence length range from hundreds to thousands, with an average document length of $231$ words. The aim of this binary sentiment classification task is to decide whether a movie review is positive or negative. Several works considered this task [248, 249, 86, 25, 233].
@@
DNA Modeling.   DNA can be easily connected to language, given that it consists of a sequence of discrete tokens (i.e., nucleotides) belonging to a limited vocabulary. Indeed, multiple LLMs solutions for genomics have been proposed. Additionally, recent works highlighted the advantages brought by considering long-range interaction when predicting gene expression [8]. Recent works in the context of this survey [203, 85] explore scaling laws across model size and sequence length as long as downstream classification. Pretraining is performed via causal language modeling (next token prediction) on the Human Genome HG38 dataset, which consists of a single human genome composed by $4.5$ billion tokens (DNA base pairs) in the training split. Please refer to the referred papers for further details.
@@ 7.4 Speech/Audio Benchmarks
Processing sequences of “raw” audio data can be challenging for neural models, due to high-frequency sampling, resulting in very long sequences. Some traditional approaches involve complex pipelines that require extracting mixed-and-matched hand-crafted features. The Speech Commands [297] dataset provides 1-second raw audio waveforms sampled at $16000$Hz (i.e., a 1-D sequence of $L=16000$), consisting of $105,829$ recordings from $2,618$ speakers – both background noise and a vocabulary of $35$ spoken words (such as ‘left’, ‘right’, etc.). The task is to classify which word was spoken. While several works [142, 131] exploit a pre-processed variant with standard mel-frequency cepstrum coefficients (MFCC), recent models tackle the problem directly processing the raw signal [90, 88, 94, 87, 271]. Other recent works [85] explits a subset of the dataset (referred to as SC09) consisting of spoken digits “zero” through “nine” ($31,158$ training utterances—$8.7$ hours in total—-by $2,032$ speakers). This is usually tackled as an autoregressive speech generation problem. Other similar dataset are: TIMIT [74], leveraged in [68]; Beethoven (with the setting described by [188]) consists of recordings of Beethoven’s 32 piano sonata; YouTubeMix is a 4 hour dataset of piano music. The latter two dataset are considered in [78].

### v2

@@ 7.3 Natural Language Benchmarks
Language Modeling.   One of the peculiar characteristic behind the success of LLMs is their ability to model the generative likelihood of word sequences, i.e., they are able to predict the probabilities of future tokens. In other words, they can build powerful Language Models for a target language, being able to capture the syntax, grammar, and semantic properties. There exist several Language Modeling dataset, and many of the works described in this survey are evaluated in Language Modeling tasks. PennTreebank (PTB) [255], WikiText-103 [256] and The Pile [257] are commonly used for evaluating model performances terms of perplexity. PTB is a language model corpus containing 1 million words. Generally, models are trained to predict the next word (a variant denoted with PTB-w, evaluated with the perplexity metric) or character (referred to as PTB-c, evaluated with bits-per-character metric). The corpus contains $42,068$ sentences ($971,657$ words, average word-length of about $4.727$ characters) of varying length (the range is from $3$ to $84$ words). The vocabulary for the character level models includes $49$ unique symbols (including one for spaces). The standard train/valid/test split consists of 929929k training words, 7373k validation words, and 8282k test words, with lower-case words, numbers replaced with token $N$, newlines replaced with <<eos>>, all other punctuation removed. The vocabulary is composed by the most frequent 1010k words, with the rest of the tokens replaced by an <<unk>> token [258]. In literature, various segmentations of the corpus into training sequences of different lengths are being used ($L=50$ for PTB-w in [31, 190], $L\in(70,300)$ for PTB-w and $L=150$ for PTB-c in [58]). This task is considered in [31, 58, 117, 162, 177, 178, 190, 22]. WikiText-103 [256] is a long-range world-level language modeling benchmark. It contains 2828k articles from English Wikipedia, with an average length of 3.63.6k tokens per article. In total, it contains 103103M words. The preprocessing is more realistic with respect to PTB (e.g., vocabulary is not limited to 1010k words, only words with a count below $3$ are discarded). It is used in [168, 25, 135, 154, 74, 84, 94, 85, 103, 114, 117, 123, 130, 99, 98, 102, 119, 143, 131, 206]. The Pile dataset [257] is an $825$ GiB English text corpus designed for training large-scale language models. It is is composed of $22$ diverse and high-quality datasets, including established Natural Language Processing datasets (PG-19, Books3, English Wikipedia, etc.) and $14$ new high-quality curated language modeling datasets derived from heterogeneous sources (PubMed, ArXiv, GitHub, Stack Exchange, etc.–see Table 1 in [257] for a full description)). Besides its impact in training large language models, The Pile can serve as a benchmark for testing cross-domain knowledge and generalization ability of language models. It is exploited in [74, 146, 75, 124, 131, 38]. Among the datasets contained in The Pile, the Project Gutenberg (PG-19) [118] consists of $28,602$ full-length books from the Project Gutenberg, published prior to $1919$, hence representing distinct writing styles with respect to other datasets based on modern text. It contains $6,966,499$ English language words. When tokenized, each PG-19 book has between 5050k-100100k tokens. This datasets has been employed to test long range modeling capabilities of the approaches proposed in [129, 128, 119]. Datasets from Arxiv and GitHub were already been introduced in [259], prior to The Pile. The ArXiv dataset is a corpus of sources of technical papers from the “Mathematics” section, downloaded via the ArXiv Bulk Data Access. The GitHub dataset consists of source code extracted from diverse GitHub repositories featuring open-source licenses. In this dataset, all files within each GitHub repository are combined to form one long document. The evaluation metric commonly used with these datasets is bits-per-token (i.e., $\text{log}_{2}(\text{perplexity})$, the lower the better). Used by [128, 129, 130]. Enwik8 [260] is a character-level dataset consisting of 100100M unprocessed bytes of the XML text dump of the English Wikipedia. Its vocabulary contains $204$ characters, and it is exploited in [81, 114, 117, 123, 164].
@@
Commonsense Reasoning.   A common experimental setting for evaluating LLMs consists in pre-training the model on The Pile and then testing the model capabilities on downstream zero-shot evaluation tasks. In particular, recent works leverage the Language Model Evaluation Harness (LM-Eval-Harness) [^25] https://github.com/EleutherAI/lm-evaluation-harness from EleutherAI [261], which mainly evaluates on tasks/datasets that measure common sense reasoning—in which the model should use “common sense” or world knowledge to perform inference. It is composed by several tasks: LAMBADA, HellaSwag, PIQA, ARC, WinoGrande. LAMBADA [262] tests the modeling capacity on long-range contextual reasoning and language comprehension abilities. The model is given a certain context sentence, and it is asked to predict the last word of a target sentence based on the context. Accuracy and perplexity of the predicted last words are measured to assess the performance on language modeling tasks. The scaling law property [263, 264] is often measured, meaning that scaling language models would improve the accuracy and reduce the perplexity. LAMBADA contains novels from the Book Corpus [265], and comprises $10,022$ passages, divided into $4,869$ development (that can be used to finetune models) and $5,153$ test passages (extracted from $1,331$ and $1,332$ disjoint novels, respectively). The average passage consists of $4.6$ sentences in the context plus $1$ target sentence, for a total length of $\approx 75$ tokens. The dataset provides also training data, which include the full text of $2,662$ novels (disjoint from those in dev/test sets), comprising $203$ million words. HellaSwag [266] [^26] Short for Harder Endings, Longer contexts, and Lowshot Activities for Situations With Adversarial Generations. expands the SWAG dataset [267], which introduced the task of commonsense natural language inference: the model receives context from a video caption, along with four potential outcomes for the subsequent events. Among the provided choices, only one is correct, representing the actual succeeding caption in the video. Negatives are created using Adversarial Filtering, a method where a set of discriminators is employed to choose a tricky collection of incorrectly generated answers. According to the HellaSwag paper, models like BERT struggle with robust commonsense reasoning and tend to learn specific biases from the dataset. If there’s a slight shift in language distribution, their performance drops significantly, even if the domain remains the same. HellaSwag introduces $70,000$ problems that are easy for humans (95.695.6% accuracy) but challenging for machines (5050%). These problems consist of video captions from the ActivityNet Captions dataset and context and follow-up paragraphs from WikiHow. To assess how a model can adapt to novel scenarios, category labels sourced from WikiHow and ActivityNet are sampled to build ’zero-shot’ evaluation sets. For each set, being it used for validation or testing, two subsets are generated. The first one includes $5,000$ ’in-domain’ examples from categories encountered during training. The second subset comprises $5,000$ ’zero-shot’ examples randomly selected from categories held out during training. In total, there are $70,000$ examples in the dataset. PIQA (Physical Interaction: Question Answering) [268] benchmarks are about physical commonsense understanding. Given a “physical” goal (“to separate egg whites using a water bottle…”) expressed in natural language and two possible solutions (($i$) “Squeeze the water bottle and press it against the yolk. Release, which creates suction and lifts the yolk;” ($ii$) “Place the water bottle and press it against the yolk. Keep pushing, which creates suction and lifts the yolk.”), a model must choose the most appropriate solution. The dataset assesses the capacity of natural language understanding models to connect text with a robust intuitive-physics model of the world. Humans effortlessly choose answer ($i$) because separating the egg involves pulling out the yolk, a task that machines find easily misleading. The dataset consists of more than $16,000$ training QA pairs with an additional $\approx 2$k and $\approx 3$k QA held out for development and testing, respectively. Goal sentences have an average length of $7.8$ words, while correct/incorrect solutions have an average length of $21.3$ words. This results in over $3.7$ million lexical tokens within the training data. The AI2 Reasoning Challenge (ARC) [269] consists of a collection of $7787$ natural grade-school science questions ( typically $4$-way multiple choice), authored for human standardized tests. It is partitioned into two sets: Challenge Set (ARC-c), which contains $2590$ questions which are incorrectly answered by both a retrieval-based algorithm and a word co-occurrence algorithm. An Easy Set (ARC-easy) contrains the remainder of the questions. WinoGrande [270] is a large-scale dataset of 4444k pronoun resolution problems. The task is to identify to which of two subject a pronoun [^27] In practice it is formatted as a fill-in-the-blank problem, where the blank corresponds to the mention of one of the two names in the context. is referring to (e.g., “Robert woke up at 9:00am while Samuel woke up at 6:00am, so _ had less time to get ready for school.”; Options: Robert / Samuel). Generally, there is a twin sentence with nearly identical structure but where the answer is the opposite. It has been designed to be unsolvable for statistical models that rely on basic word associations. Going back to the main container of the just described datasets, the LM-Eval-Harness, we mentioned that it has been used in several recent works [146, 44, 145, 38, 75, 141]. Apart from LM-Eval-Harness, other comprehensive benchmarks for commons sense reasoning have been tested. Measuring Massive Multitask Language Understanding (MMLU) [271] is a versatile benchmark for large-scale evaluation of multi-task knowledge understanding. It consists of $57$ tasks including elementary mathematics, US history, computer science, law, etc. It collects $15,908$ questions in total, split into a few-shot development set, a validation set, and a test set. As shown in existing work, LLMs mostly outperform small models by a substantial margin on this benchmark, also showing scaling laws in model size. We point the reader attention towards a recent survey on LLMs [272] for further details on other commonly used datasets, such as CMMLU [273], C-Eval [274] BoolQ [275], OpenBookQA [276] and others.
@@
Machine Translation.   Another sequence-oriented benchmark in the context of language generation is conditional text generation [277], which focuses on generating texts satisfying specific task demands based on the given conditions. A common task is machine translation [278], whose goal is the translation of text or speech from a source language to another. Among the most common benchmarks, the WMT (several editions—2014 [279] up to 2022 [280]) contains translation tasks from/to English to/from several languages, such as Czech, French, German, Hindi, and Russian. A common testbed is to use the WMT’14 English-to-German, which contains 4.6M sentence pairs, and the WMT’17 the Chinese-to-English split which is composed by 20.620.6M sentence pairs. The case-sensitive NIST BLEU score [281] is commonly adopted as the evaluation metric. It has been tackled in [135, 84, 85, 112, 113, 99, 110]. A similar benchmark is the IWSLT German–English spoken-domain translation [282]. It consists of $209,772$ sentence pairs from transcribed TED and TEDx presentations, with a mean sentence length of $103$ characters for German and $93$ for English. It has been considered in [84, 162].
@@
Other Language-oriented Benchmarks.   Apart from the generation of high-quality natural language text, several models show strong abilities to generate regular and formal languages, such as computer programs (i.e., code synthesis) [114, 272]. Another popular direction for evaluation is the one is sentiment classification. The principal testbed is the IMDb movie review dataset [283], a collection of 5050k movie reviews, equally balanced into positive and negative reviews and processed into equal-size train and test sets. Sequence length range from hundreds to thousands, with an average document length of $231$ words. The aim of this binary sentiment classification task is to decide whether a movie review is positive or negative. Several works considered this task [184, 186, 39, 162, 102].
@@
DNA Modeling.  DNA can be easily connected to language, given that it consists of a sequence of discrete tokens (i.e., nucleotides) belonging to a limited vocabulary. Indeed, multiple LLM-based solutions for genomics have been proposed. Additionally, recent works have highlighted the advantages of by considering long-range interaction when predicting gene expression [284]. Recent works in the context of this survey [157, 146] explore scaling laws across model size and sequence length as well as downstream classification. Pretraining is performed via causal language modeling (next token prediction) on the Human Genome HG38 dataset, which consists of a single human genome containing $4.5$ billion tokens (DNA base pairs) in the training split. Please refer to the cited papers for further details.
@@ 7.4 Speech/Audio Benchmarks
Processing sequences of “raw” audio data can be challenging for neural models, due to high-frequency sampling, resulting in very long sequences. Some traditional approaches involve complex pipelines that require extracting mixed-and-matched hand-crafted features. The Speech Commands [285] dataset provides 1-second raw audio waveforms sampled at 1600016000Hz (i.e., a 1-D sequence of $L=16000$), consisting of $105,829$ recordings from $2,618$ speakers – both background noise and a vocabulary of $35$ spoken words (such as ‘left’, ‘right’, etc.). The task is to classify which word was spoken. While several works [286, 178] exploit a pre-processed variant with standard mel-frequency cepstrum coefficients (MFCC), recent models tackle the problem directly processing the raw signal [40, 25, 149, 150, 42]. Other recent works [146] explits a subset of the dataset (referred to as SC09) consisting of spoken digits “zero” through “nine” ($31,158$ training utterances—$8.7$ hours in total—-by $2,032$ speakers). This is usually tackled as an autoregressive speech generation problem. Other similar dataset are: TIMIT [287], leveraged in [175]; Beethoven (with the setting described by [288]) consists of recordings of Beethoven’s 32 piano sonata; YouTubeMix is a 4 hour dataset of piano music. The latter two dataset are considered in [148].